OpenStudy (anonymous):

What is the percent decrease of 100,000 to 25,000. Show your work.

OpenStudy (kira_yamato):

25000/100000 = 0.25 0.25 = 25%

OpenStudy (anonymous):

that is how you get percent decrease by dividing

OpenStudy (anonymous):

\[\frac{100000-25000}{100000}100= 75\% \]